OpenStudy (anonymous):

Suppose A and B are both divisible by M. Prove that A^2-B^2 is divisible by M.

OpenStudy (perl):

A is divisible by M : A= M*k, k is some integer B is divisible by M: B = M*h , h is some integer Now A^2 - B^2 = (M*k)^2 - (M*h)^2 = M^2*k^2 - M^2*h^2 = M^2 * (k^2 - h^2) = M * [M(k^2-h^2)] so A^2 - B^2 = M*(some integer) therefore A^2 - B^2 is divisible by M

OpenStudy (perl):

you can make a stronger statement, A^2 - B^2 is divisible by M^2

OpenStudy (anonymous):

or you could say if A is divisible by M and B is divisible by M then also \(A-B\) is divisible by M and \(A+B\) is as well